# Temperature-dependent Raman spectra of 2H-MoS2 from Machine Learning-driven statistical sampling

### 02 Apr 2026 v1

@@ Abstract
Molybdenum sulfides are in the spotlight of materials science thanks to their interesting properties for applications in optoelectronics, nanocomposites, lubricants, and catalysis. The structural characterization of Molybdenum sulfides is a crucial step to understand and tune their properties. Vibrational techniques, such as infrared and Raman spectroscopy, can directly link to structural features, but the experimental literature suffers from large variability. Theoretical calculations are a powerful tool complementing and explaining empirical measurements. The reliability of first-principles calculation depends on the level of approximation made, taking into account disorder, doping, or temperature to yield a good description of the phonon statistics and related measurable quantities, such as the infrared and Raman peaks.
In this study we calculate the Raman spectrum of crystalline 2H-MoS2, including broadening and shifts due to thermal and anharmonic effects. Our results demonstrate excellent agreement with experimental measurements; notably, the calculated temperature trends in frequencies and linewidths align with empirical observations. These findings establish a robust computational framework, paving the way for similar studies on amorphous Molybdenum sulfides.
@@ I Introduction
Transition Metal Chalcogenides (TMCs) are promising materials for applications in optoelectronics [75, 44], catalysis [28, 25, 30, 16], as lubricants [67, 59], in the polymer industry [51, 71, 1], and even for elastic energy storage [53]. Notably, Molybdenum Sulfides (MSs) exhibit promising catalytic activity in the Hydrogen Evolution Reaction (HER) [25, 38, 76], representing a non-toxic, cheap and less polluting alternative to current catalysts which are mostly based on Platinum and Lead [49, 10, 40]. The tailoring of amorphous MSs catalysts, which show the highest activity [4, 41], is still strongly hindered by uncertainty on the actual electrolytic mechanism [52, 65]. This is also due to the absence of long-range order in these materials, which makes characterizing their structure a very hard challenge.
Raman spectroscopy is among the cheapest and most accessible characterization techniques, and is widely used to study TMCs [74, 55, 42]. As an example, in layered systems, the peak frequencies are extremely sensitive to the interlayer interaction, enabling quick identification of systems which are mono, bi-layer, etc… [34]. Experimentally many other effects can also lead to peak shifts, from strain and substrate epitaxy to dielectric environment and defects.
@@
An approximation to the Raman spectrum can also be calculated at the harmonic level, with a few additional dielectric calculations. The quasi-harmonic approximation [2] can be used to account for the thermal expansion, but does account for peak shifts due to explicit anharmonicity [73]. Furthermore, the shape of the Raman peaks, caused by the finite phonon lifetime and anharmonicity, is usually accounted for as an empirical broadening of a Lorentzian bell. To account for both the thermal expansion and the anharmonicity, we rely on the Temperature Dependent Effective Potential [56, 60, 32, 24, 23, 31] (TDEP) method. Another element which often is missing in 0 K calculations is the quantum zero-point renormalization (ZPR); we investigate its effect on by comparing two different sampling methods: classical MD (no ZPR) and a stochastic method using quantum occupations (with ZPR). As TDEP itself requires a representative and accurate sampling of forces and positions (see II.3), we leverage the Moment Tensor Potential [50] (MTP) Machine Learning Interatomic Potential (MLIP) to achieve accuracies close to DFT with a fraction of the cost.
The rest of this paper is structured as follows: in Section II we present the computational methods that were used throughout the work. We start from the description of the MLIP and its training and testing procedure, followed by the sampling methods. Then we address the calculation of the vibrational properties within the TDEP framework, and, finally, the calculation of the Raman quantities. In Section III we report and discuss our results. First we assess the performance of our MLIP, then we move to the determination of the unit cells at finite temperature. Next we investigate the phonon band dispersions and the spectral function properties, as compared to experiments. We conclude this Section presenting the Raman results. In the last Section, IV, we recapitulate and briefly present some future perspectives.
@@ II.1 The MLIP
The ideal dataset used to train the MLIP should be representative of the structures the MLIP is meant compute the properties of in production. In our case, we are interested in exploring the canonical phase space of bulk hexagonal MoS2 between 100 and 700 K. For MTP (as for the majority of MLIPs), the only input to the model (except for the hyperparameters) is the chemical identity and the position of each atom in the system; therefore, we need realistic atomic positions in the above-mentioned temperature range, accounting for anharmonicity and phonon occupation. To obtain a collection of such structures (or configurations) we employed the Machine Learning Assisted Canonical Sampling algorithm [8, 9] (MLACS). The key idea is that starting from an initial structure (the 2H-MoS2 crystal structure from Materials Project database [27], id: mp-2815) the ground state (GS) properties, such as energy, atomic forces and stress tensor, are computed via DFT (with Abinit [68, 57, 19, 66]) and used to perform a first training of the MLIP. With ML-driven Molecular Dynamics (MLMD) a short time evolution ($\sim$100 fs) (or trajectory) are simulated, to explore the phase space. The GS properties and structures of selected configurations along the trajectory are stored in the dataset and the steps are repeated from the MLIP training. Thanks to the iterative dataset augmentation, the MLIP improves its description of the phase space regions already visited and therefore the MD simulation progressively gains in accuracy. To adequately cover the temperature-dependent extent of the phase space, at each MLACS iteration, a random temperature in the range is selected at every MLACS iteration for the MD simulation. It is crucial, for the first MLACS iterations, to keep the simulations short, because the MLIP suffers from two quality-limiting factors: overfitting due to the small size of the dataset, and the early stage of the iterative procedure. As the MLIP gets better, it is safe to set longer simulation times. To decide when to stop the loop, a convergence assessment strategy can be employed. We consider MLACS to be converged when the incorporation of new configurations into the dataset does not increase the fit error of the model trained on it, meaning that the new additions lie in a region already known by the MLIP. This is done by performing a k-fold cross-validation every N new configurations. For each dataset size, the error of the cross-validation is calculated as the average over the k folds, and used as our convergence criterion.
@@ II.2.2 Molecular dynamics sampling
Note that for both sampling methods, the equilibrium geometry at finite temperature must be defined, either by experiments (e.g. temperature-dependent XRD), numerically as NPT thermal averages, or via the quasi-harmonic approximation method. Our equilibrium cells are obtained as time averages of 40 independent NPT MD simulations with the same starting point, each lasting 95 ps (after 5 ps of equilibration, discarded) and sampled every 50 fs. To assess the convergence with respect to the number of instances, we compute differences of the $a$ and $c$ cell parameters obtained with 1, 2, 3 and 4 instances less. Then we do the same calculation for many shuffles of the instances and averaged the result. At any temperature, with 40 iterations we obtain average absolute differences below $5\cdot 10^{-6}$Å. Finally, in both cases, the supercell is defined by expanding the hexagonal unit cell via a matrix of the shape:
@@ II.3 Extraction of the vibrational properties
$U(\mathbf{R-\mathbf{R}}_{eq})=U(\mathbf{u})=U(0)+\sum_{ij\alpha\beta}\left.\frac{1}{2}\frac{\partial^{2}U}{\partial u_{i\alpha}\partial u_{j\beta}}\right|_{0}u_{i\alpha}u_{j\beta}+\sum_{ijk\alpha\beta\gamma}\left.\frac{1}{2}\frac{\partial^{3}U}{\partial u_{i\alpha}\partial u_{j\beta}\partial u_{k\gamma}}\right|_{0}u_{i\alpha}u_{j\beta}u_{k\gamma}+$...
$=U(0)+\sum_{i\alpha}u_{i\alpha}\underbrace{\left[\sum_{j\beta}\frac{1}{2}\Phi_{i\alpha j\beta}^{(2)}\cdot u_{j\beta}\right]}_{\equiv F_{i\alpha}^{(2)}}+\sum_{i\alpha}u_{i\alpha}\underbrace{\sum_{jk\beta\gamma}\left[\frac{1}{2}\Phi_{i\alpha j\beta k\gamma}^{(3)}\cdot u_{j\beta}u_{k\gamma}\right]}_{\equiv F_{i\alpha}^{(3)}}+$... (1)
where $x_{i\alpha}$ is the displacement of atom $i$ along direction $\alpha$, $\Phi_{i\alpha j\beta}^{(2)}$ is the IFC2 between atom $i$ along $\alpha$ and atom $j$ along $\beta$, $\Phi_{i\alpha j\beta k\gamma}^{(3)}$ is the third-order IFC (IFC3). We also define the two-body force acting on atom $i$ along $\alpha$ due to the displacement of atom $j$ as $F_{i\alpha}$ (and similarly for the 3-body forces). Note that a first guess of the IFCs allows to initialize the model, and yields a tentative calculation of the atomic forces. The IFC2s are obtained by minimizing the average difference between the reference forces and those computed with the model. The higher-order IFCs are calculated sequentially on the residual of the previous orders. For example, the IFC3s are obtained by minimizing the error in the residual forces after removing the second-order part.
The first-order term in equation (II.3) is zero at the equilibrium geometry by construction. However, the finite-temperature NVT dynamics can produce differences in the average atomic positions. In this case, the NPT cell (alone) with the 0 K reduced coordinates will not yield the equilibrium geometry, and the perturbative approach does not hold, leading to non-zero first-order terms. To overcome this problem, we fit the first-order terms as well and displace (via gradient descent) the atoms towards new positions. By repeating this operation iteratively, the atomic reference positions converge to a minimum-energy geometry [24].
@@
Here $Z^{*}_{i\alpha\gamma}=\partial U^{2}/\partial u_{i\alpha}\partial\mathcal{E}_{\gamma}$ is the Born-effective charge associated to displacement of atom $i$ along $\alpha$, producing an electrical polarization along $\gamma$, . $q_{\gamma}$ is the $\gamma$ component of the phonon wavevector $\mathbf{q}$ and $\varepsilon^{\infty}_{\gamma\delta}$ is the dielectric susceptibility along $\delta$ due to an electric field along $\gamma$.
@@
while $c_{i}^{\kappa}$ and $M_{i}^{\kappa}$ are, respectively, the fractional concentration and mass of isotope $\kappa$ of atom $i$, and $\bar{M}_{i}$, and $\bar{M}_{i}=\sum_{\kappa}c_{i}^{\kappa}m_{i}^{\kappa}$.
@@
where $\Delta(\omega)$ and $\Gamma(\omega)$ are, respectively, the real and imaginary parts of the self-energy. Note that according to this definition the self-energy has units of squared frequency, but TDEP follows the convention of normalizing it by a factor of $2\omega_{\lambda}$: $\Delta_{\lambda}(\omega)^{\mathrm{TDEP}}\propto\Delta_{\lambda}(\omega)/\omega_{\lambda}$ (and similarly for $\Gamma$). The imaginary part of the self-energy is computed from the IFC3s as [56]:
@@
which represents a sum over all three-phonon scattering events with $\Phi_{\lambda\lambda^{\prime}\lambda^{\prime\prime}}$ being the IFC3s in phonon space, and $n_{\lambda}$ being the Bose-Einstein occupation. The real part is obtained via Kramers-Kronig transformations. We converge the error on the FWHMs below 0.005 THz for the 64x64x64 summing grid, with respect to the 56x56x56 case. The choice of the resolution of the frequency axis is also critical, especially when the spectral function is used to compute non-integrated quantities (such as the Raman spectrum). In such cases, the local shape of the peaks is determinant and it can be dramatically affected by low resolution. For our spectral functions, we converge to a resolution below 0.0034 THz. Note that sTDEP and MD-TDEP method are justified
@@ II.4 Calculation of the Raman spectrum
where $\omega_{e}$ is the frequency of excitation, $\omega$ is the Raman shift, $\omega_{\lambda}$ is the harmonic frequency of the phonon, and $\mathbf{\hat{e}}_{s}$ and $\mathbf{\hat{e}}_{i}$ are the polarization vectors of the incoming and scattered light, respectively. With $n(\omega_{\lambda},T)$ we denote the occupation of the phonon mode at temperature $T$ (excluding the zero-point-motion), while $\tilde{A}_{\lambda}(\omega)$ is the spectral function, normalized to $1$. Finally, $\partial\chi^{\prime}/\partial Q_{\lambda}$ is the intensive Raman tensor:
@@
It is important to specify that (7) is valid for eigenvectors of the harmonic Hamiltonian defined with the position operator $Q_{j\gamma}\propto 1/\sqrt{2\omega_{\lambda}}(a^{\dagger}_{-\mathbf{q}_{\lambda}}+a_{\mathbf{q}_{\lambda}})$, which produces the $1/\omega_{\lambda}$ factor in the equation. Note that according to (7) the spectrum given by the normalized spectral function is multiplied by a constant of the Raman shift (but its value depends on the phonon frequency), if we consider $\omega_{e}-\omega\simeq\omega_{e}$.
@@ II.5 Note on polarization
$\alpha_{\lambda}^{\prime}$ $=$ $\frac{1}{3}\tilde{\alpha}^{\prime}_{xx}+\tilde{\alpha}^{\prime}_{yy}+\tilde{\alpha}^{\prime}_{zz}$

### 09 Apr 2026 v2

@@ Abstract
Molybdenum sulfides are in the spotlight of materials science thanks to their interesting properties for applications in optoelectronics, nanocomposites, lubricants, and catalysis. The structural characterization of Molybdenum sulfides is a crucial step to understand and tune their properties. Vibrational techniques, such as infrared and Raman spectroscopy, can directly link to structural features, but the experimental literature suffers from large variability. Theoretical calculations are a powerful tool complementing and explaining empirical measurements. The reliability of first-principles calculation depends on the level of approximation made, taking into account disorder, doping, or temperature to yield a good description of the phonon statistics and related measurable quantities, such as the infrared and Raman peaks. In this study we calculate the Raman spectrum of crystalline 2H-MoS2, including broadening and shifts due to thermal and anharmonic effects. Our results demonstrate excellent agreement with experimental measurements; notably, the calculated temperature trends in frequencies and linewidths align with empirical observations. These findings establish a robust computational framework, paving the way for similar studies on amorphous Molybdenum sulfides.
@@ I Introduction
Transition Metal Chalcogenides (TMCs) are promising materials for applications in optoelectronics [75, 44], catalysis [28, 25, 30, 16], as lubricants [67, 59], in the polymer industry [51, 71, 1], and even for elastic energy storage [53]. Notably, Molybdenum Sulfides (MSs) exhibit promising catalytic activity in the Hydrogen evolution reaction [25, 38, 76], representing a non-toxic, cheap and less polluting alternative to current catalysts which are mostly based on Platinum and Lead [49, 10, 40]. The tailoring of amorphous MSs catalysts, which show the highest activity [4, 41], is still strongly hindered by uncertainty on the actual electrolytic mechanism [52, 65]. This is also due to the absence of long-range order in these materials, which makes characterizing their structure a very hard challenge.
Raman spectroscopy is among the cheapest and most accessible characterization techniques, and is widely used to study TMCs [74, 55, 42]. As an example, in layered systems, the peak frequencies are extremely sensitive to the interlayer interaction, enabling quick identification of systems which are mono-, bi-layer, etc… [34]. Experimentally many other effects can also lead to peak shifts, from strain and substrate epitaxy to dielectric environment and defects.
@@
An approximation to the Raman spectrum can also be calculated at the harmonic level with a few additional dielectric calculations. The quasi-harmonic approximation [2] can be used to incorporate thermal expansion, but does not account for peak shifts due to explicit anharmonicity [73]. Furthermore, the shape of the Raman peaks, caused by the finite phonon lifetime and anharmonicity, is usually described as an empirical broadening of a Lorentzian bell. To capture both thermal expansion and anharmonicity, we rely on the Temperature Dependent Effective Potential [56, 60, 32, 24, 23, 31] (TDEP) method. Another element which is often missing in 0 K calculations is the quantum zero-point renormalization (ZPR); we investigate its effect by comparing two different sampling methods: classical MD (no ZPR) and a stochastic method using quantum occupations (with ZPR). As TDEP itself requires a representative and accurate sampling of forces and positions (see II.3), we leverage the Moment Tensor Potential [50] (MTP) Machine Learning Interatomic Potential (MLIP) to achieve accuracies close to DFT with a fraction of the cost.
The rest of this paper is structured as follows: in Section II we present the computational methods that were used throughout the work. We start with the description of the MLIP and its training and testing procedure, followed by the sampling methods. Then we address the calculation of the vibrational properties within the TDEP framework, and, finally, the calculation of the Raman quantities. In Section III we report and discuss our results. First we assess the performance of our MLIP, then we move to the determination of the unit cells at finite temperature. Next we investigate the phonon band dispersions and the spectral function properties, as compared to experiments. We conclude this Section presenting the Raman results. In the last Section, IV, we recapitulate and briefly present some future perspectives.
@@ II.1 The MLIP
The ideal dataset used to train the MLIP should be representative of the structures the MLIP is meant to compute the properties of in production. In our case, we are interested in exploring the canonical phase space of bulk hexagonal MoS2 between 100 and 700 K. For MTP (as for the majority of MLIPs), the only input to the model (except for the hyperparameters) is the chemical identity and the position of each atom in the system; therefore, we need realistic atomic positions in the above-mentioned temperature range, accounting for anharmonicity and phonon occupation. To obtain a collection of such structures (or configurations) we employed the Machine Learning Assisted Canonical Sampling algorithm [8, 9] (MLACS). The key idea is that starting from an initial structure (the 2H-MoS2 crystal structure from Materials Project database [27], id: mp-2815) the ground state (GS) properties, such as energy, atomic forces and stress tensor, are computed via DFT (with Abinit [68, 57, 19, 66]) and used to perform a first training of the MLIP. With ML-driven Molecular Dynamics (MLMD) a short time evolution ($\sim$100 fs) (or trajectory) is simulated, to explore the phase space. The GS properties and structures of selected configurations along the trajectory are stored in the dataset and the steps are repeated from the MLIP training. Thanks to the iterative dataset augmentation, the MLIP improves its description of the phase space regions already visited and therefore the MD simulation progressively gains in accuracy. To adequately cover the temperature-dependent extent of the phase space, at each MLACS iteration, a random temperature in the range is selected at every MLACS iteration for the MD simulation. It is crucial, for the first MLACS iterations, to keep the simulations short, because the MLIP suffers from two quality-limiting factors: overfitting due to the small size of the dataset, and the early stage of the iterative procedure. As the MLIP gets better, it is safe to set longer simulation times. To decide when to stop the loop, a convergence assessment strategy can be employed. We consider MLACS to be converged when the incorporation of new configurations into the dataset does not increase the fit error of the model trained on it, meaning that the new additions lie in a region already known by the MLIP. This is done by performing a k-fold cross-validation every N new configurations. For each dataset size, the error of the cross-validation is calculated as the average over the k folds, and used as our convergence criterion.
@@ II.2.2 Molecular dynamics sampling
Note that for both sampling methods, the equilibrium geometry at finite temperature must be defined, either by experiments (e.g. temperature-dependent XRD), numerically as NPT thermal averages, or via the quasi-harmonic approximation method. Our equilibrium cells are obtained as time averages of 40 independent NPT MD simulations with the same starting point, each lasting 95 ps (after 5 ps of equilibration, discarded) and sampled every 50 fs. To assess the convergence with respect to the number of instances, we compute differences of the $a$ and $c$ cell parameters obtained with 1, 2, 3 and 4 instances less. Then we do the same calculation for many shuffles of the instances and average the result. At any temperature, with 40 iterations we obtain average absolute differences below $5\cdot 10^{-6}$Å. Finally, in both cases, the supercell is defined by expanding the hexagonal unit cell via a matrix of the shape:
@@ II.3 Extraction of the vibrational properties
$U(\mathbf{R-\mathbf{R}}_{eq})=U(\mathbf{u})=U(0)+\sum_{ij\alpha\beta}\left.\frac{1}{2}\frac{\partial^{2}U}{\partial u_{i\alpha}\partial u_{j\beta}}\right|_{0}u_{i\alpha}u_{j\beta}+\sum_{ijk\alpha\beta\gamma}\left.\frac{1}{6}\frac{\partial^{3}U}{\partial u_{i\alpha}\partial u_{j\beta}\partial u_{k\gamma}}\right|_{0}u_{i\alpha}u_{j\beta}u_{k\gamma}+$...
$=U(0)+\sum_{i\alpha}u_{i\alpha}\underbrace{\left[\sum_{j\beta}\frac{1}{2}\Phi_{i\alpha j\beta}^{(2)}\cdot u_{j\beta}\right]}_{\equiv F_{i\alpha}^{(2)}}+\sum_{i\alpha}u_{i\alpha}\underbrace{\sum_{jk\beta\gamma}\left[\frac{1}{6}\Phi_{i\alpha j\beta k\gamma}^{(3)}\cdot u_{j\beta}u_{k\gamma}\right]}_{\equiv F_{i\alpha}^{(3)}}+$... (1)
where $u_{i\alpha}$ is the displacement of atom $i$ along direction $\alpha$, $\Phi_{i\alpha j\beta}^{(2)}$ is the IFC2 between atom $i$ along $\alpha$ and atom $j$ along $\beta$, $\Phi_{i\alpha j\beta k\gamma}^{(3)}$ is the third-order IFC (IFC3). We also define the two-body force acting on atom $i$ along $\alpha$ due to the displacement of the other atoms as $F_{i\alpha}^{(2)}$ (and similarly for the 3-body forces). Note that a first guess of the IFCs allows to initialize the model, and yields a tentative calculation of the atomic forces. The IFC2s are obtained by minimizing the average difference between the reference forces and those computed with the model. The higher-order IFCs are calculated sequentially on the residual of the previous orders. For example, the IFC3s are obtained by minimizing the error in the residual forces after removing the second-order part.
The first-order terms in equation (II.3) (not shown) are zero at the equilibrium geometry by construction. However, the finite-temperature NVT dynamics can produce differences in the average atomic positions. In this case, the NPT cell (alone) with the 0 K reduced coordinates will not yield the equilibrium geometry, and the perturbative approach does not hold, leading to non-zero first-order terms. To overcome this problem, we fit the first-order terms as well and displace (via gradient descent) the atoms towards new positions. By repeating this operation iteratively, the atomic reference positions converge to a minimum-energy geometry [24].
@@
Here $Z^{*}_{i\alpha\gamma}=\partial U^{2}/\partial u_{i\alpha}\partial\mathcal{E}_{\gamma}$ is the Born-effective charge associated to displacement of atom $i$ along $\alpha$, producing an electrical polarization along $\gamma$. $q_{\gamma}$ is the $\gamma$ component of the phonon wavevector $\mathbf{q}$ and $\varepsilon^{\infty}_{\gamma\delta}$ is the dielectric susceptibility along $\delta$ due to an electric field along $\gamma$.
@@
while $c_{i}^{\kappa}$ and $M_{i}^{\kappa}$ are, respectively, the fractional concentration and mass of isotope $\kappa$ of atom $i$, and $\bar{M}_{i}=\sum_{\kappa}c_{i}^{\kappa}m_{i}^{\kappa}$.
@@
where $\Delta_{\lambda}(\omega)$ and $\Gamma_{\lambda}(\omega)$ are, respectively, the real and imaginary parts of the self-energy. Note that according to this definition the self-energy has units of squared frequency, but TDEP follows the convention of normalizing it by a factor of $2\omega_{\lambda}$: $\Delta_{\lambda}(\omega)^{\mathrm{TDEP}}\propto\Delta_{\lambda}(\omega)/\omega_{\lambda}$ (and similarly for $\Gamma$). The imaginary part of the self-energy is computed from the IFC3s as [56]:
@@
which represents a sum over all three-phonon scattering events with $\Phi_{\lambda\lambda^{\prime}\lambda^{\prime\prime}}$ being the IFC3s in phonon space, and $n_{\lambda}$ being the Bose-Einstein occupation. The real part is obtained via Kramers-Kronig transformations. We converge the error on the FWHMs below 0.005 THz for the 64x64x64 summing grid, with respect to the 56x56x56 case. The choice of the resolution of the frequency axis is also critical, especially when the spectral function is used to compute non-integrated quantities (such as the Raman spectrum). In such cases, the local shape of the peaks is determinant and it can be dramatically affected by low resolution. For our spectral functions, we converge to a resolution below 0.0034 THz.
@@ II.4 Calculation of the Raman spectrum
where $\omega_{e}$ is the frequency of excitation, $\omega$ is the Raman shift, $\omega_{\lambda}$ is the harmonic frequency of the phonon, and $\mathbf{\hat{e}}_{s}$ and $\mathbf{\hat{e}}_{i}$ are the polarization vectors of the incoming and scattered light, respectively. With $n(\omega_{\lambda},T)$ we denote the occupation of the phonon mode at temperature $T$ (excluding the zero-point-motion), while $\tilde{A}_{\lambda}(\omega)$ is the spectral function, normalized to 11. Finally, $\partial\chi^{\prime}/\partial Q_{\lambda}$ is the intensive Raman tensor:
@@
It is important to specify that (7) is valid for eigenvectors of the harmonic Hamiltonian defined with the position operator $Q_{\lambda}\propto 1/\sqrt{2\omega_{\lambda}}(a^{\dagger}_{-\mathbf{q}_{\lambda}}+a_{\mathbf{q}_{\lambda}})$, which produces the $1/\omega_{\lambda}$ factor in the equation. Note that according to (7) the spectrum given by the normalized spectral function is multiplied by a constant of the Raman shift (but its value depends on the phonon frequency), if we consider $\omega_{e}-\omega\simeq\omega_{e}$.
@@ II.5 Note on polarization
$\alpha^{\prime}$ $=$ $\frac{1}{3}\tilde{\alpha}^{\prime}_{xx}+\tilde{\alpha}^{\prime}_{yy}+\tilde{\alpha}^{\prime}_{zz}$